OpenStudy (anonymous):

Write the equation of the line containing the points (5, 5) and (-5, 0). Use the form y = mx + b. A. y=1/2x+5/2 B. y=2x+10 C. y=-1/2x-5/2 D. y=-2x-10

OpenStudy (dmndlife24):

We first need to find the slope containing those two points... Formula: \[m=\frac{ y2-y1 }{ x2-x1 }\] If our points are (5, 5) and (-5, 0), then... y2 = 0 y1 = 5 x2 = -5 x1 = 5

OpenStudy (dmndlife24):

\[\frac{ (0) - (5) }{ (-5) - (5) }\]

OpenStudy (dmndlife24):

\[m = \frac{ -5 }{ -10 } = \frac{ 1 }{ 2 }\]

OpenStudy (dmndlife24):

From finding the slope, you should already know the answer :)

OpenStudy (anonymous):

yes thank you soo much ah i just really needed a refresher i learned all of this at the beginning of the year :)))